# Disentangling the development of collective flow in high energy proton proton collisions with a multiphase transport model

### v1

@@ Abstract
In this work, we investigate the collective flow development in high energy proton proton (pp) collisions with a multiphase transport model (AMPT) based on PYTHIA8 initial conditions with a sub-nucleon structure. It is found that the PYTHIA8 based AMPT model can reasonably describe both the charged hadron productions and elliptic flow experimental data measured in pp collisions at $\sqrt{s}=13$ TeV. By turning on the parton and hadron rescatterings in AMPT separately, we find that the observed collective flow in pp collisions is largely developed during the parton evolutions, while no significant flow effect can be generated with the pure hadronic rescatterings. It is also shown that the parton escape mechanism is important for describing both the magnitude of the two-particle cumulant and the sign of the four-particle cumulants. We emphasize that the strong mass ordering of the elliptic flow results from the coalescence process in the transport model can thus be regarded as unique evidence related to the creation of deconfined parton matter in high energy pp collisions.
@@ 1 Introduction
In the past decade, collective flow like signals have also been surprisingly observed in numerous experimental data for small systems created in the high multiplicity proton nucleus (pA) collisions, proton proton (pp) collisions and even photon nucleus collisions which were not expected to flow Khachatryan:2010gv ; Abelev:2012ola ; Khachatryan:2016txc ; Adare:2014keg ; Khachatryan:2015waa ; PHENIX:2018lia ; Acharya:2019vdf ; ATLAS:2021jhn . On aside of the continuing experimental efforts to further examine the resemblance between the small systems and the conventional large systems made by nucleus nucleus (AA) collisions, multiple theoretical frameworks have been constructed to understand the origin of the collective flow effect in small systems. One school of the theoretical models assumes no contribution of deconfined parton matter evolution in small systems and employs either initial state correlations Schenke:2014zha ; Schenke:2016lrs ; Zhang:2020ayy ; Altinoluk:2020psk ; Shi:2020djm or final state interactions between overlapped strings and cascading hadrons Bierlich:2015rha ; Bierlich:2016vgw ; Bierlich:2017vhg ; Bierlich:2020naj ; Bierlich:2021poz ; Zhou:2015iba to generate the collectivity. The other school of the phenomenological studies incorporate a parton evolution phase dominated by the quark and gluon degrees of freedom, in which hydrodynamic expansions or parton cascade effects lead to the sizable azimuthal anisotropies via the final state interactions in the parton evolution stage He:2015hfa ; Deng:2011at ; Ma:2014pva ; Zhao:2017rgg ; Zhao:2020pty ; Zhao:2022ayk ; Zhao:2021bef ; Greif:2017bnr ; YuanyuanWang:2023meu . With all the hot debates on the different mechanisms to induce the collective flow in small systems, it is extremely important to further investigate the collective flow in terms of its multi-particle correlations and $p_{\rm T}$ dependence in a systematic way. In this work, we focus on the understanding of the flow development in high energy pp collisions.
@@ 2 The model and analysis method
Figure 1: The pseudo-rapidity distribution (a) and the transverse momentum spectra (b) for charged hadrons in pp collisions at $\sqrt{s}=13$ TeV from the AMPT model calculations (solid line) compared to ALICE data(black circle) with at least one charged hadron in the mid-rapidity region (INEL$>$0 events). ALICE data are taken from Ref. Adam:2015pza
Figure 2: Average transverse momentum for $\pi$, K and proton within $|y|<0.5$ and $30<p_{\rm T}<$ 3 GeV$/c$ varying with the charged particle density from the AMPT model calculations (solid lines) compared to ALICE data(open and solid markers) in pp collisions at $\sqrt{s}=13$ TeV. ALICE data are taken from Ref. Acharya:2020zji .
@@ 3.1 Hadron productions
To begin with, we compare the charged particle productions in pp collisions at $\sqrt{s}=13$ TeV of the AMPT model calculations with PYTHIA8 initial conditions to the experimental data. The results of the model calculations are obtained when all the final state parton/hadron rescatterings and resonance decays are finished. Events with at least one charged track in mid-rapidity $|\eta|<1$ are selected for the comparison (INEL$>$0 events). It is found in Fig. 1(a) that the charged particle density distributions per unit pseudorapidity space measured in the experiments Adam:2015pza are well described by the model calculations. We also present the results of charged hadron transverse momentum spectra in Fig. 1(b). The model calculations are found to reproduce the experimental data of transverse momentum spectra in a reasonable way.
@@
We explore the event activity dependence of identified hadron productions by examining the average transverse momentum $\langle p_{\rm T}\rangle$ for $\pi^{\pm}$, $K^{\pm}$, $p$ and $\bar{p}$ versus the charged particle multiplicity in Fig. 2. The identified particles within $|y|<0.5$ having $30<p_{\rm T}<$ 3 GeV$/c$ are used in this analysis. The results are obtained following the experimental procedure in Ref. Acharya:2020zji . In order to restrict the comparison to the low $p_{\rm T}$ regime where the string melting version of the AMPT model is more applicable, the experimental data are extracted by refitting the measured $p_{\rm T}$ spectra using the Tsallis distribution in every event class within $30<p_{\rm T}<$ 3 GeV$/c$. The experimentally observed strong increase of average $p_{\rm T}$ with the increasing event multiplicity is reproduced in the model calculations. The mass hierarchy of $\langle p_{\rm T}\rangle$, which is connected to the radial flow effect, is found to be naturally included in this framework. It is shown in our previous study Zheng:2021jrr that this multiplicity dependence of the average $p_{\rm T}$ for different hadrons arises from the convolution of multiple parton interactions implemented in the PYTHIA8 initial conditions together with the final state interactions. Compared to our previous results in Ref. Zheng:2021jrr , the parton scattering cross section has been tuned from 0.2 mb to 0.15mb in this work to have a better agreement with the elliptic flow data observed in pp collisions at $\sqrt{s}=13$ TeV. It is also seen that the hadron yield and transverse momentum distributions are generally not sensitive to the parton scattering cross sections.
@@
Figure 3: Two-particle cumulants in the sub-event method (a) and four-particle cumulants (b) for charged hadrons with $|\eta|<2.4$ and $30.3<p_{\rm T}<$ 3 GeV$/c$ in pp collisions at $\sqrt{s}=13$ TeV versus the average selected charged hadron number $\langle N_{ch}\rangle$ from the model calculations and the experimental data. The dash dotted line in (b) represents results obtained with event averaging performed directly with $N_{ch}$. CMS data are taken from Ref. CMS:2016fnw . ATLAS data are taken from Ref. ATLAS:2017hap .
@@
Figure 4: Two-particle cumulant (a) and extracted $v_{2}$ coefficients in high multiplicity events with $M>100$ (b) based on the sub-event method using $|\Delta\eta|>2$ after subtracting the non-flow contribution with low multiplicity events for charged hadrons in pp collisions at $\sqrt{s}=13$ TeV. Charged hadrons with $|\eta|<2.4$ and $30.3<p_{\rm T}<$ 3 GeV$/c$ are defined as reference flow particles in this analysis. CMS data are taken from Ref. CMS:2016fnw . ATLAS data are taken from Ref. ATLAS:2016yzd .
@@ 3.2 Azimuthal flow with Q-cumulant method
We present the two-particle cumulants $c_{2}\{2\}$ after applying $\Delta\eta$ cut with the sub-event method and four-particle cumulants $c_{2}\{4\}$ for charged hadrons in Fig. 3. The impacts of final state parton and hadron evolution stages on the building of multi-particle cumulants are explored by switching on each key ingredient implemented in AMPT step by step. In Fig. 3(a), the black open circles and red open squares represent the $c_{2}\{2\}$ experimental data with $|\Delta\eta|>2$ measured by CMS CMS:2016fnw and ATLAS ATLAS:2017hap , respectively. The red curve shows the results from AMPT calculations without any final state parton or hadron rescatterings. The green curve and the blue curve represents the scenario with only partonic interactions and hadronic interactions, respectively. The result with all the final state interactions is shown by the gray curve. The bands represent the statistical uncertainties of the model calculations. It is found that with all the final state parton and hadron interactions included, the model calculation describes the observed charged two-particle cumulants reasonably, except at $\langle N_{ch}\rangle$ below 30. Comparing the partonic and hadronic final state interactions, the sizable $c_{2}\{2\}$ cumulants appear mainly during the parton evolution stage and the inclusion of additional hadronic rescattering generally diminishes the two particle correlations . With only hadronic rescattering effect, the result is quite close to the case without any final state interactions, compatible with the long range correlation studies Zheng:2021jrr . It is also shown that the results either with or without final state interactions become quite similar in the region of $\langle N_{ch}\rangle$ less than 20, indicating non-flow effects are dominant in the low multiplicity region.
@@
Multi-particle cumulant $c_{2}\{4\}$ is supposed to be less sensitive to the non-flow effects and uncover the collective motion of the particle system in a more exclusive way. It is expected for AA collisions that $c_{2}\{4\}$ should be negative if the correlation comes from the hydrodynamic collective flow Bzdak:2013rya ; Zhao:2021bef ; Jia:2017hbm . In Fig. 3(b), we show the comparisons for the four-particle cumulants $c_{2}\{4\}$ varying with the event multiplicity represented by average charged particle number $\langle N_{ch}\rangle$. The discrepancy between ATLAS and CMS data arises due to the treatment on the selection of the events. Events are analyzed in reference particle number bins and then mapped to corresponding $N_{ch}$ values in ATLAS data while event averaging is performed directly with $N_{ch}$ in CMS data ATLAS:2017hap . We find that the observed negative $c_{2}\{4\}$ in high multiplicity events from AMPT is predominantly coming from the parton evolution stage similar to the arise of the sizable two-particle cumulant as seen in Fig. 3(a). This is similar to the finding of Ref. Zhao:2024wqs , which shows that the partonic interactions are essential for the flow like correlations in high multiplicity jets. $c_{2}\{4\}$ with only final state parton rescatterings reaches smaller value in the negative region. The inset shows that with only partonic interactions $c_{2}\{4\}$ becomes negative at similar $N_{ch}$ compared to the scenario that both partonic and hadronic final interactions are turned on. With all the final state parton and hadron interactions, the gray band follows the decreasing trend as the increase of the event multiplicity, qualitatively similar to the behavior of ATLAS data. It is also noticed that the gray band turns into negative earlier than the ATLAS data. By using an event averaging method based on the number of selected charged particles similar to the CMS analysis method represented by the dash-dotted line in Fig. 3(b), we confirm that the model predictions with all the final state effects become close to the CMS data, reinstating the discrepancy between ATLAS and CMS data in our calculations.
Figure 5: The $p_{\rm T}$ differential $v_{2}$ (left column) and $v_{2}^{sub}$ (right column) of identified hadrons based on the sub-event method using pseudorapidity gap $|\Delta\eta|>2$ for high multiplicity events with $M>100$ in pp collisions at $\sqrt{s}=13$ TeV. The AMPT results with all the final state interactions, only parton rescatterings, only hadronic rescatterings and no final state interactions are shown in the first to the fourth row, respectively.
@@
To further investigate the flow coefficients with the two-particle correlation method, we subtract the non-flow contribution using the two-particle cumulants in low multiplicity events from $c_{2}\{2\}$ in the sub-event method following the prescription in Eq. 12. It is observed in Fig. 4(a) that the subtracted $c_{2}\{2\}$ is approaching zero in the low multiplicity region for all model setups, suggesting the impact of non-flow subtraction. However, one can see that the non-flow effect is not completely removed as the red line is not at zero at finite $N_{ch}$. The final $c_{2}\{2\}$ with both parton and hadron rescatterings is initially growing with $N_{ch}$ and then saturated following the trend of the experimental data. The ATLAS data are slightly higher than the CMS data in Fig. 4(a) simply because a higher $p_{T}$ cut $50.5<p_{\rm T}<$ 5 GeV$/c$ has been applied for charged hadron selections in that analysis ATLAS:2016yzd .It is also found that with only hadronic final state interactions, no collective flow signal can be generated except in the very high multiplicity region. In Fig. 4(b), we investigate the $p_{\rm T}$ differential charged particle $v_{2}$ after low multiplicity non-flow subtraction for the high multiplicity events with $M>100$. The subtracted $p_{\rm T}$ differential $v_{2}$ data are found to be well described by the model calculations at final state. $v_{2}^{sub}\{2\}(p_{T})$ is increasing with the increase of $p_{\rm T}$ and becomes saturated in the higher $p_{\rm T}$ region. Without the partonic evolution stage (shown by the red and blue curves), the subtracted $v_{2}$ is close to zero across the whole $p_{\rm T}$ region.
@@
After implementing the low multiplicity subtraction method deduced in Sect. 2, the mass ordering feature of $v_{2}^{sub}$ in each model setup is similar to the $v_{2}$ case as is shown in the right column of Fig. 5. Although the $v_{2}^{sub}$ magnitude decreases after subtractions, the result indicates that the mass ordering effect uncovered in the comparison is robust either with or without the non-flow contaminations. It is also observed that the subtracted $v_{2}^{sub}$ for all hadron species become close to zero in the case with only hadronic rescatterings or without any final state interactions, making it impossible to identify any mass ordering features in these two scenarios. This observation suggests that the sizable $v_{2}$ shown in Fig. 5 (e) and (g) are therefore dominant by the non-flow contributions.

### v2

@@ Abstract
In this work, we investigate the collective flow development in high energy proton proton (pp) collisions with a multiphase transport model (AMPT) based on PYTHIA8 initial conditions with a sub-nucleon structure. It is found that the PYTHIA8 based AMPT model can reasonably describe both the charged hadron productions and elliptic flow experimental data measured in pp collisions at $\sqrt{s}=13$ TeV. By turning on the parton and hadron rescatterings in AMPT separately, we find that the observed collective flow in pp collisions is largely developed during the parton evolution, while no significant flow effect can be generated with the pure hadronic rescatterings. It is also shown that the parton escape mechanism is important for describing both the magnitude of the two-particle cumulant and the sign of the four-particle cumulants. We emphasize that the strong mass ordering of the elliptic flow results from the coalescence process in the transport model and can thus be regarded as unique evidence related to the creation of deconfined parton matter in high energy pp collisions.
@@ 1 Introduction
In the past decade, collective flow like signals have also been surprisingly observed in numerous experimental data for small systems created in the high multiplicity proton nucleus (pA) collisions, proton proton (pp) collisions and even photon nucleus collisions which were not expected to flow Khachatryan:2010gv ; Abelev:2012ola ; Khachatryan:2016txc ; Adare:2014keg ; Khachatryan:2015waa ; PHENIX:2018lia ; Acharya:2019vdf ; ATLAS:2021jhn . On aside of the continuing experimental efforts to further examine the resemblance between the small systems and the conventional large systems made by nucleus nucleus (AA) collisions, multiple theoretical frameworks have been constructed to understand the origin of the collective flow effect in small systems. One school of the theoretical models assumes no contribution of deconfined parton matter evolution in small systems and employs either initial state correlations Schenke:2014zha ; Schenke:2016lrs ; Zhang:2020ayy ; Altinoluk:2020psk ; Shi:2020djm or final state interactions among partons dEnterria:2010xip , overlapped strings Bierlich:2015rha ; Bierlich:2016vgw ; Bierlich:2017vhg ; Bierlich:2020naj and cascading hadrons Bierlich:2021poz ; Zhou:2015iba to generate the collectivity. The other school of the phenomenological studies incorporate a parton evolution phase dominated by the quark and gluon degrees of freedom, in which hydrodynamic expansions or parton cascade effects lead to the sizable azimuthal anisotropies via the final state interactions in the parton evolution stage He:2015hfa ; Deng:2011at ; Ma:2014pva ; Zhao:2017rgg ; Zhao:2020pty ; Zhao:2022ayk ; Zhao:2021bef ; Greif:2017bnr ; YuanyuanWang:2023meu . With all the hot debates on the different mechanisms to induce the collective flow in small systems, it is extremely important to further investigate the collective flow in terms of its multi-particle correlations and $p_{\rm T}$ dependence in a systematic way. In this work, we focus on the understanding of the flow development in high energy pp collisions.
@@ 2 The model and analysis method
Figure 1: (Colour Online) The pseudo-rapidity distribution (a) and the transverse momentum spectra (b) for charged hadrons in pp collisions at $\sqrt{s}=13$ TeV from the AMPT model calculations with all final state parton and hadron rescatterings (allFSI) and without final state interactions (noFSI) compared to ALICE data with at least one charged hadron in the mid-rapidity region (INEL$>$0 events). The allFSI and noFSI (only $dN/d\eta$) results are represented by the solid lines and dashed lines, respectively. ALICE data indicated by the open circles are taken from Ref. Adam:2015pza
Figure 2: (Colour Online) Average transverse momentum for $\pi$, K and proton within $|y|<0.5$ and $30<p_{\rm T}<$ 3 GeV$/c$ varying with the charged particle density from the AMPT model calculations with all final state parton and hadron rescatterings (allFSI) and without final state interactions (noFSI) compared to ALICE data in pp collisions at $\sqrt{s}=13$ TeV. The allFSI and noFSI results are represented by the solid lines and dashed lines, respectively. ALICE data indicated by the markers are taken from Ref. Acharya:2020zji .
@@ 3.1 Hadron productions
To begin with, we compare the charged particle productions in pp collisions at $\sqrt{s}=13$ TeV of the AMPT model calculations with PYTHIA8 initial conditions to the experimental data. The results of the model calculations shown by the solid lines are obtained when all the final state parton/hadron rescatterings and resonance decays are finished. Events with at least one charged track in mid-rapidity $|\eta|<1$ are selected for the comparison (INEL$>$0 events). It is found in Fig. 1(a) that the charged particle density distributions per unit pseudorapidity space measured in the experiments Adam:2015pza are well described by the model calculations. We also present the results of charged hadron transverse momentum spectra in Fig. 1(b). The model calculations are found to reproduce the experimental data of transverse momentum spectra in a reasonable way. We also present the results without any final state interactions in Fig. 1 shown by the red dashed line. The results without any final state interactions are supposed to be similar to the pure PYTHIA8 predictions. It can be found in this comparison that the final state interactions and coalescence hadronization mechanism incorporated in the AMPT framework will slightly reduce the multiplicity number at final state for high energy pp collisions.
@@
We explore the event activity dependence of identified hadron productions by examining the average transverse momentum $\langle p_{\rm T}\rangle$ for $\pi^{\pm}$, $K^{\pm}$, $p$ and $\bar{p}$ versus the charged particle multiplicity in Fig. 2. The identified particles within $|y|<0.5$ having $30<p_{\rm T}<$ 3 GeV$/c$ are used in this analysis. The results are obtained following the experimental procedure in Ref. Acharya:2020zji . In order to restrict the comparison to the low $p_{\rm T}$ regime where the string melting version of the AMPT model is more applicable, the experimental data are extracted by refitting the measured $p_{\rm T}$ spectra using the Tsallis distribution in every event class within $30<p_{\rm T}<$ 3 GeV$/c$. The experimentally observed strong increase of average $p_{\rm T}$ with the increasing event multiplicity is reproduced in the model calculations. The mass hierarchy of $\langle p_{\rm T}\rangle$, which is connected to the radial flow effect, is found to be naturally included in this framework. The results with and without final state interactions are presented by the solid and dashed lines in this figure. The observation is consistent with our previous study Zheng:2021jrr that the multiplicity dependence of the average $p_{\rm T}$ for different hadrons arises from the convolution of multiple parton interactions implemented in the PYTHIA8 initial conditions. The final state interactions further enlarge the mass hierarchy of the $p_{\rm T}$ distributions for different particle species and result in the slight overestimation of the average $p_{\rm T}$ for protons. Compared to our previous results in Ref. Zheng:2021jrr , the parton scattering cross section has been tuned from 0.2 mb to 0.15mb in this work to have a better agreement with the elliptic flow data observed in pp collisions at $\sqrt{s}=13$ TeV. It is also seen that the hadron yield and transverse momentum distributions are generally not sensitive to the parton scattering cross sections.
@@
Figure 3: (Colour Online) Two-particle cumulants in the sub-event method (a) and four-particle cumulants (b) for charged hadrons with $|\eta|<2.4$ and $30.3<p_{\rm T}<$ 3 GeV$/c$ in pp collisions at $\sqrt{s}=13$ TeV versus the average selected charged hadron number $\langle N_{ch}\rangle$ from the model calculations and the experimental data. The gray solid lines represent the calculations with all final state interactions (0.15mb with hFSI). The results with only final state parton rescatterings (0.15mb no hFSI), only final state hadron rescatterings (0mb with hFSI) and no final state interactions (no FSI) are shown by the green dash-dotted lines, blue dotted lines and red dashed lines, respectively. The bands represent the statistical uncertainties. CMS data shown by the open circles are taken from Ref. CMS:2016fnw . ATLAS data shown by the open squares are taken from Ref. ATLAS:2017hap .
@@
Figure 4: (Colour Online) Two-particle cumulant (a) and extracted $v_{2}$ coefficients in high multiplicity events with $M>100$ (b) based on the sub-event method using $|\Delta\eta|>2$ after subtracting the non-flow contribution with low multiplicity events for charged hadrons in pp collisions at $\sqrt{s}=13$ TeV. Charged hadrons with $|\eta|<2.4$ and $30.3<p_{\rm T}<$ 3 GeV$/c$ are defined as reference flow particles in this analysis. The gray solid lines represent the calculations with all final state interactions (0.15mb with hFSI). The results with only final state parton rescatterings (0.15mb no hFSI), only final state hadron rescatterings (0mb with hFSI) and no final state interactions (no FSI) are shown by the green dash-dotted lines, blue dotted lines and red dashed lines, respectively. The bands represent the statistical uncertainties. CMS data shown by the open circles are taken from Ref. CMS:2016fnw . ATLAS data shown by the open squares are taken from Ref. ATLAS:2016yzd .
Figure 5: (Colour Online) The $p_{\rm T}$ differential $v_{2}$ (left column) and $v_{2}^{sub}$ (right column) of identified hadrons based on the sub-event method using pseudorapidity gap $|\Delta\eta|>2$ for high multiplicity events with $M>100$ in pp collisions at $\sqrt{s}=13$ TeV. The AMPT results with all the final state interactions (0.15mb with hFSI), only parton rescatterings (0.15mb no hFSI), only hadronic rescatterings (0mb with hFSI) and no final state interactions (noFSI) are shown in the first to the fourth row, respectively. The $\pi^{\pm}$, $K^{\pm}$, and proton results are shown by gray solid lines, blue dotted lines and red dashed lines, respectively. The bands represent the statistical uncertainties.
@@ 3.2 Azimuthal flow with Q-cumulant method
We present the two-particle cumulants $c_{2}\{2\}$ after applying $\Delta\eta$ cut with the sub-event method and four-particle cumulants $c_{2}\{4\}$ for charged hadrons in Fig. 3. The impacts of final state parton and hadron evolution stages on the building of multi-particle cumulants are explored by switching on each key ingredient implemented in AMPT step by step. In Fig. 3(a), the black open circles and red open squares represent the $c_{2}\{2\}$ experimental data with $|\Delta\eta|>2$ measured by CMS CMS:2016fnw and ATLAS ATLAS:2017hap , respectively. The red dashed line shows the results from AMPT calculations without any final state parton or hadron rescatterings. The green dash-dotted line and the blue dotted line represents the scenario with only partonic interactions and hadronic interactions, respectively. The result with all the final state interactions is shown by the gray solid line. The bands represent the statistical uncertainties of the model calculations. It is found that with all the final state parton and hadron interactions included, the model calculation describes the observed charged two-particle cumulants reasonably, except at $\langle N_{ch}\rangle$ below 30. Comparing the partonic and hadronic final state interactions, the sizable $c_{2}\{2\}$ cumulants appear mainly during the parton evolution stage and the inclusion of additional hadronic rescattering generally diminishes the two particle correlations. The reduction of $c_{2}\{2\}$ after incorporating hadronic final state interactions might be related to the rapid isotropization of the parton matter created in pp collisions. The sizable initial eccentricity induced through sub-nucleon spatial fluctuations can drop rapidly during the expansion of the parton system. The follow-up hadronic scatterings will decrease the momentum space anisotropy with a more isotropized spatial configuration. With only hadronic rescattering effect, the result is quite close to the case without any final state interactions, compatible with the long range correlation studies Zheng:2021jrr . It is also shown that the results either with or without final state interactions become quite similar in the region of $\langle N_{ch}\rangle$ less than 20, indicating non-flow effects are dominant in the low multiplicity region.
@@
Multi-particle cumulant $c_{2}\{4\}$ is supposed to be less sensitive to the non-flow effects and uncover the collective motion of the particle system in a more exclusive way. It is expected for AA collisions that $c_{2}\{4\}$ should be negative if the correlation comes from the hydrodynamic collective flow Bzdak:2013rya ; Zhao:2021bef ; Jia:2017hbm . In Fig. 3(b), we show the comparisons for the four-particle cumulants $c_{2}\{4\}$ varying with the event multiplicity represented by average charged particle number $\langle N_{ch}\rangle$. The discrepancy between ATLAS and CMS data arises due to the treatment on the selection of the events. Events are analyzed in reference particle number bins and then mapped to corresponding $N_{ch}$ values in ATLAS data while event averaging is performed directly with $N_{ch}$ in CMS data ATLAS:2017hap . We find that the observed negative $c_{2}\{4\}$ in high multiplicity events from AMPT is predominantly coming from the parton evolution stage similar to the arise of the sizable two-particle cumulant as seen in Fig. 3(a). This is similar to the finding of Ref. Zhao:2024wqs , which shows that the partonic interactions are essential for the flow like correlations in high multiplicity jets. $c_{2}\{4\}$ with only final state parton rescatterings reaches smaller value in the negative region. The inset shows that with only partonic interactions $c_{2}\{4\}$ becomes negative at similar $N_{ch}$ compared to the scenario that both partonic and hadronic final interactions are turned on. With all the final state parton and hadron interactions, the gray line follows the decreasing trend as the increase of the event multiplicity, qualitatively similar to the behavior of ATLAS data. It is also noticed that the gray line turns into negative earlier than the ATLAS data. We also confirmed the event averaging effect with our model calculations. By using an event averaging method based on the number of selected charged particles similar to the CMS analysis procedure, the results with all the final state effects become close to the CMS data, reinstating the discrepancy between ATLAS and CMS data.
@@
To further investigate the flow coefficients with the two-particle correlation method, we subtract the non-flow contribution using the two-particle cumulants in low multiplicity events from $c_{2}\{2\}$ in the sub-event method following the prescription in Eq. 12. It is observed in Fig. 4(a) that the subtracted $c_{2}\{2\}$ is approaching zero in the low multiplicity region for all model setups, suggesting the impact of non-flow subtraction. However, one can see that the non-flow effect is not completely removed as the red line is not at zero at finite $N_{ch}$. The final $c_{2}\{2\}$ with both parton and hadron rescatterings is initially growing with $N_{ch}$ and then saturated following the trend of the experimental data. The ATLAS data are slightly higher than the CMS data in Fig. 4(a) simply because a higher $p_{T}$ cut $50.5<p_{\rm T}<$ 5 GeV$/c$ has been applied for charged hadron selections in that analysis ATLAS:2016yzd . It is also found that with only hadronic final state interactions, no collective flow signal can be generated except in the very high multiplicity region. In Fig. 4(b), we investigate the $p_{\rm T}$ differential charged particle $v_{2}$ after low multiplicity non-flow subtraction for the high multiplicity events with $M>100$. The subtracted $p_{\rm T}$ differential $v_{2}$ data are found to be well described by the model calculations at final state. $v_{2}^{sub}\{2\}(p_{T})$ is increasing with the increase of $p_{\rm T}$ and becomes saturated in the higher $p_{\rm T}$ region. Without the partonic evolution stage (shown by the red dashed lines and the blue dotted lines), the subtracted $v_{2}$ is close to zero across the whole $p_{\rm T}$ region.
@@
After implementing the low multiplicity subtraction method deduced in Sect. 2, the mass ordering feature of $v_{2}^{sub}$ in each model setup is similar to the $v_{2}$ case as is shown in the right column of Fig. 5. Although the $v_{2}^{sub}$ magnitude decreases after subtractions, the result indicates that the mass ordering effect uncovered in the comparison is robust either with or without the non-flow contaminations. It is also observed that the subtracted $v_{2}^{sub}$ for all hadron species become close to zero in the case with only hadronic rescatterings or without any final state interactions, making it impossible to identify any mass ordering features in these two scenarios. This observation suggests that the sizable $v_{2}$ shown in Fig. 5 (e) and (g) are therefore dominant by the non-flow contributions. In the comparisons of this section, we find that the final state parton evolution phase and the coalescence hadronization process coupled to it are important ingredients to produce the sizable elliptic flow signal and significant mass ordering features in high energy pp collisions. It has been shown that very few collisions in transport model approach might generate significant flow feature in small systems like pp collisions He:2015hfa ; Greif:2017bnr ; Kurkela:2018ygx ; Borghini:2010hy . The observations in this work can be another example suggesting that the full hydrodynamization process may not be necessary to explain the experimentally observed collectivity in pp collisions Grosse-Oetringhaus:2024bwr .
Figure 6: (Colour Online) Two-particle cumulants in the sub-event method after the low multiplicity events non-flow contribution is subtracted(a) and four-particle cumulants (b) for final state partons with $|\eta|<2.4$ and $30.3<p_{T}<$ 3 GeV$/c$ in pp collisions at $\sqrt{s}=13$ TeV versus the number of reference partons in the flow calculations. Gray solid lines represent the normal calculations while red dashed lines represent the results with random $\phi$ test. The bands represent the statistical uncertainties.
Figure 7: (Colour Online) The $p_{\rm T}$ differential $v_{2}$ (a) and $v_{2}^{sub}$ (b) of identified hadrons with random $\phi$ test based on the sub-event method using pseudorapidity gap $|\Delta\eta|>2$ for high multiplicity events with $M>100$ in pp collisions at $\sqrt{s}=13$ TeV. The $\pi^{\pm}$, $K^{\pm}$, and proton results are shown by gray solid lines, blue dotted lines and red dashed lines, respectively. The bands represent the statistical uncertainties.